OpenStudy (anonymous):

What is the parametric equation for 9 =(x+3)^2+(y-1)^2????

OpenStudy (anonymous):

this is circle of radius 3, so lets use these substitutions for x and y: x = 3cost y = 3sint plugging that in we get: \[9 = (3\cos(t)+3)^{2} +(3\sin(t)-1)^{2}\] Simplifying gives us: \[9 = 9\cos^{2}(t)+18\cos(t)+9+9\sin^{2}(t)-6\sin(t)+1\] \[9 = 19+18\cos(t)-6\sin(t) \Rightarrow -10 = 18\cos(t)-6\sin(t) \Rightarrow -5 = 9\cos(t)-3\sin(t)\]

OpenStudy (anonymous):

how did you get a radius of 3 ?

OpenStudy (anonymous):

equation of a circle is: \[(x-h)^{2}+(y-k)^{2} = r^{2}\] (h,k) is the center of the circle, r is the radius. Thus: \[r^{2} = 9 \Rightarrow r = 3\]

OpenStudy (anonymous):

If you need anything else explained let me know :)

OpenStudy (anonymous):

gotcha......th

OpenStudy (anonymous):

gotcha......th

OpenStudy (anonymous):

gotcha......th

OpenStudy (anonymous):

gotcha......th

OpenStudy (anonymous):

your simplifying is a little hard to understand...

OpenStudy (anonymous):

sry bout that, one key thing is: \[9\cos^{2}+9\sin^{2} = 9(\cos^{2}+\sin^{2}) = 9(1) = 9\] ( took out the t cause im lazy >.< imagine its there >.>)